OpenStudy (anonymous):

the measures of the angles of a triangle are in the ration 1:4:7. what are the measures of the angles?

OpenStudy (anikay):

Well, we know that all triangles add up to 180 degrees right?

OpenStudy (anonymous):

yes

OpenStudy (anikay):

so can you set up an equation given your ratios and the knowledge that they all add up to 180?

OpenStudy (anonymous):

need just an example to know how ..to follow..not sure how to set up the equation and I remember I need to solve x

OpenStudy (anikay):

x+4x+7x=180

OpenStudy (anikay):

do you understand how I did that?

OpenStudy (anonymous):

I did get this far just not sure where to go ..the numbers are representing theangles but there is where I get lost.

OpenStudy (anikay):

you then combine like terms\[12x = 180\]

OpenStudy (anikay):

because 1+4+7=12

OpenStudy (anonymous):

ok that looks easy

OpenStudy (anonymous):

x+4x+7x = 180 12+48x+84x=180 ??

OpenStudy (anikay):

um

OpenStudy (anikay):

lol sorry, no

OpenStudy (anonymous):

not sure where to go and how to get the 180..

OpenStudy (anikay):

the first equation is correct. and if you simplify it you get 12x=180

OpenStudy (anikay):

then how do you find x?

OpenStudy (anonymous):

not sure

OpenStudy (anikay):

algebra mannnn \[x=\frac{ 180 }{ 12 }\] x= 15

OpenStudy (anikay):

right?

OpenStudy (anonymous):

thxs i can no do the remain problems

OpenStudy (anikay):

but, we're not done yet, the problem asks for all the angles

OpenStudy (anikay):

and since x=15 that's just the first the ratio is 1:4:7, so what are the rest?

OpenStudy (anonymous):

2:6:11

OpenStudy (anonymous):

x =21/180 x =8.57 2(8.57)+6(8.57)+11(8.57) 17.14 +51.42 + 94.27 = 180

OpenStudy (anikay):

have the right idea sir

OpenStudy (anikay):

however, you added wrong. 2+6+11=19 180/19 = 9.474 the rest of what you did was right